The dynamic plate resistance of a triode valve is 10 kΩ. Find the change in the plate current if the plate voltage is changed from 200 V to 220 V.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. Dynamic Plate Resistance r base p = 10 KΩ = 10^4 Ω δI base p = ? δV base p = 220 – 220 = 20 V δI base p = (δV base p / r base p) / V base g = constant. = 20/10^4 = 0.002 A = 2 mA